OpenStudy (anonymous):

Factor and simplify the algebraic expression. (x+7)^1/4 - (x+7)^5/4

OpenStudy (anonymous):

\[(x+7)^{\frac{1}{4}}-(x+7)^{\frac{5}{4}}\] \[=(x+7)^{\frac{1}{4}}(1-(x+7)^{\frac{4}{4}})\] is a start

OpenStudy (anonymous):

Isn't (x+7)^1/4 the GCF?

OpenStudy (anonymous):

then then next step, since \(\frac{4}{4}=1\) is to rewrite as \[=(x+7)^{\frac{1}{4}}(1-(x+7))\]

OpenStudy (anonymous):

it is the GCF and that is why you factor it out

OpenStudy (anonymous):

so my original equation = what you just gave me?

OpenStudy (anonymous):

the "final answer" would be \[(x+7)^{\frac{1}{4}}(-x-6)\] or \[-(x+6)(x+7)^{\frac{1}{4}}\] sicnce \(1-(x+7)=-x-6\)

OpenStudy (anonymous):

But I thought that (1-(x+7)) = -x-7 because you distribute the negative and then the one makes it stay the same... right?